x(y^2+z)-y(x^2+z)q=(x^2-y^2)z, x+y=0, z=1


1
Expert's answer
2021-02-24T07:28:19-0500

"\frac{dx}{x(y^2+z)}=\frac{dy}{-y(x^2+z)}=\frac{dz}{z(x^2-y^2)}"

"\frac{xdx+ydy-dz}{x^2y^2+x^2z-x^2y^2-zy^2-zx^2+zy^2}=\frac{xdx+ydy-dz}{0}"

"xdx+ydy-dz=0"

"\frac{x^2}{2}+\frac{y^2}{2}-z=c_1"


"\frac{dx/x+dy/y+dz/z}{y^2+z-x^2-z+x^2-y^2}=\frac{dx/x+dy/y+dz/z}{0}"

"lnx+lny+lnz=lnc_2"

"xyz=c_2"